1. A refrigerator and a rice cooker are connected in parallel to a 220 V line. The current in the refrigerator is 10 A and in the rice cooker is 4 A. Find the following;



(a) the resistance of each device


(b) the total resistance


(c) the current flowing in the line

1
Expert's answer
2022-04-15T16:40:08-0400

a) The resistance of refrigerator:


"R_r=\frac{U}{I_r}=\frac{220}{10}=22\space Ohm."

The resistance of rice coocker:

"R_c=\frac{U}{I_c}=\frac{220}{4}=55\space Ohm."


b) The total resistance:


"R=\frac{R_rR_c}{R_r+R_c}=\frac{22\cdot 55}{22+55}=15.7\space Ohm."


c) The current flowing in the line:


"I=I_r+I_c=10+4=14\space A."